# Understanding the Challenges of Maker Entrepreneurship

### v1

@@ 6. Conclusion
Our interview study with 20 maker entrepreneurs, six creative entrepreneurs, and seven people who support creative entrepreneurs highlights the tension these makers feel between their creative passions and the practical challenges of entrepreneurship. Their needs go beyond that of other similarly-sized small businesses, in large part because of the way that the creative aspects of their work affect their identity and values. These differences have important ramifications for technologies and training to support this group.
Our research highlights a need for more tailored solutions in the context of maker entrepreneurs’ small-scale, physical creative goods-oriented businesses. Based on our findings, we provide recommendations for designers to consider in their development of tools to provide business support for maker entrepreneurs, focusing on supporting the unique identity, values, and training needs of maker entrepreneurs. Overall, we intend for this work to contribute to the intersection of HCI and entrepreneurship and to the development of new digital tools to help maker entrepreneurs operate successful businesses.

### v2

@@ 6. Conclusion
Our interview study with 20 maker entrepreneurs, six creative entrepreneurs, and seven people who support creative entrepreneurs
highlights the tension these makers feel between their creative passions and the practical challenges of entrepreneurship. Their needs go beyond that of other similarly-sized small businesses, in large part because of the way that the creative aspects of their work affect their identity and values. These differences have important ramifications for technologies and training to support this group. Our research highlights a need for more tailored solutions in the context of maker entrepreneurs’ small-scale, physical creative goods-oriented businesses. Based on our findings, we provide recommendations for designers to consider in their development of tools to provide business support for maker entrepreneurs, focusing on supporting the unique identity, values, and training needs of maker entrepreneurs. Overall, we intend for this work to contribute to the intersection of HCI and entrepreneurship and to the development of new digital tools to help maker entrepreneurs operate successful businesses.